(a) The graph of y=f(x) is shown. Translate it to get the graph of y=f(x+4).
(b) The graph of y=g(x) is shown. Translate it to get the graph of y=g(x)+3.

Answer by ikleyn(52925)

.
(a) The graph of y=f(x) is shown. Translate it to get the graph of y=f(x+4).
(b) The graph of y=g(x) is shown. Translate it to get the graph of y=g(x)+3.
~~~~~~~~~~~~~~~~~~~~~

(a)  To get the graph y = f(x+4), translate the graph y = f(x) horizontally 
     4 units to the left.


(a)  To get the graph y = g(x)+3, translate the graph y = g(x) vertically 
     3 units up.

Solved.